solve the system of equations by graphing
-3x-y=3
7x+y=7

Answer by sridhar(2)

-3x - y = 3
+7x + y = 7
------------
4x = 10
------------
x = 10/4 = 2.5
7x + y = 7
y = 7 - 7x
y = 7 - (7 * 2.5)
y = 7 - 17.5 = -10.5
S0, x = 2.5
y = -10.5